# A discrete Benamou-Brenier formulation of Optimal Transport on graphs

### v1

@@ Abstract
We propose a discrete transport equation on graphs which connects distributions on both vertices and edges. We then derive a discrete analogue of the Benamou-Brenier formulation for Wasserstein-$1$ distance on a graph and as a result classify all $W_{1}$ geodesics on graphs.
@@ I Introduction
Let $\mu,\nu$ be a measures on $\mathcal{X}$ and $d$ a metric on $\mathcal{X}$, we define the set of couplings of $\mu$ and $\nu$ to be $\Pi(\mu,\nu):=\{\pi\in\mathcal{P}(\mathcal{X}^{2}):\pi(A\times\mathcal{X})=\mu(A),\pi(\mathcal{X}\times B)=\nu(B)\}$. Then we define the Wasserstein $p$ distance by:
$W_{p}(\mu,\nu)^{q}=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathcal{X}^{2}}d(x,y)^{p}d\pi\right).$ (1)
It is known that $W_{p}$ is a metric on the space of (finite $p$’th moment) measures on $\mathcal{X}$, see [20, Definition 6.4]. The Wasserstein distance is popular in machine learning circles, namely as a loss function [2, 8, 17] and utilising its Riemannian structure for Wasserstein Gradient Flows [10, 15]. One notable development connecting optimal transport and differential geometry came via the Benamou-Brenier formulation, which re-parametrised the minimisation in terms of time-dependent distribution and velocity field pairs $(f_{t},v_{t})$.
@@
where $(f_{t},v_{t})$ are constrained by the transport equation - also known as the continuity equation:
$\partial_{t}f_{t}+\nabla\cdot(f_{t}v_{t})=0$ (3)
@@
The infimum in (2) is achieved by choosing $f_{t}$ as a constant speed $W_{p}$ geodesic and $v_{t}$ satisfying $v_{t}||=|\dot{f}|_{W_{p}}$, (see [5]).
On $\mathbb{R}^{d}$, the constant speed $W_{p}$ geodesics are the McCann displacement interpolations, see [14, Definition 1.1]. Though for discrete measures, say on $\{0,1...,n\}\subset\mathbb{R}$, the McCann interpolations will traverse $\mathbb{R}$, but not remain in the domain $\{0,1...,n\}$. This means to interpolate across discrete domains, considering them as embedded in a continuum will not suffice. This unfortunately sacrifices the classical Benamou-Brenier formula.
Work has already been done to amend this. Maas [13], gave a discrete transport equation and a new metric $\mathcal{W}$, distinct from $W_{2}$, which minimised a kinetic energy functional and interplayed well with discrete gradient flows.
Beckmann [4] provided a (time independent) velocity field formulation of $W_{1}$ on both continuous and discrete domains which preceded Benamou and Brenier. A concise reference can be found in [18, Section 6.3].
Finally, Hillion and Johnson [9] provided conditions for a discrete Benamou-Brenier formulation on $\mathbb{Z}$, which we are building on. In this paper we offer our own discrete transport equation including triples $(f,v,g)$, and derive a Benamou-Brenier formulation for $W_{1}$ on trees (Lemma 2, Theorem 1). We then generalise this result to graphs (Lemma 3, Theorem 2) and classify the constant speed $W_{1}$ geodesics via the solutions to our formulation (Proposition 4).
@@ II Notation and the Discrete Transport Equation
Consider the path of Binomial distributions $f(t)=\text{Bin}(n,p(t))$ on the set $\mathcal{V}=\{0,1,...,n\}$ and $p$ varies with time. We show the case $n=5$ below:
@@
where $v(t)=np^{\prime}$ and $g(t)=\text{Bin}(n-1,p)$. Our key insight is that $\mathcal{V}$ can be considered as the vertices of a path graph $\mathcal{G}$, and that $g$ and $v$ exists on the edges of $\mathcal{G}$ as follows:
@@ II-A Matrices, Operators, Equations and Integrals
As shorthand for $\omega_{x,k}=1$ and $\omega_{k,x}=-1$ we will write $x\rightarrow k$ and $k\rightarrow x$ respectively.
@@
If we consider $\nabla_{1}$ as instead defined on edges (x,x1)(x,x-1) of $\mathbb{Z}$, it corresponds exactly to the $\nabla\cdot$ operator. The composition $\mathcal{L}\coloneqq\Omega\Omega^{T}$ is the Laplacian.
@@ II-B The Discrete Transport Equation
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with incidence matrix $\Omega$, we say the transport equation is satisfied if: for a path of distributions on vertices $f=f(t)_{x}$, a path of functions on edges $v=v(t)_{k}$ (called the velocity), and a path of distributions on edges $g=g(t)_{k}$, we have
$\partial_{t}f(t)_{x}+(\nabla\cdot v(t)g(t))_{x}\text{ for all }x\in\mathcal{V},t\in[0,1].$ (4)
@@
and define $f$ to be $f(t)=Z(1,s,s,s)$ for an arbitrary decreasing and positive $s(t)$ and partition function $Z(t)=\frac{1}{1+3s}$. This is the stationary distribution of a Markov jump process $Q$ where $q_{x0}s(t)=q_{0x}$. Equation (4) states that
@@ III-A Benamou-Brenier on the Vertices of a Tree
It is worth noting that we have other special cases of $W_{1}$. For cyclic graphs, we have a another closed form expression which utilise cuts, see [19] and for a general graph, $W_{1}$ can be expressed as the minimum transport cost across possible spanning trees of $\mathcal{G}$, see [16].
@@
Let $\mathcal{G}$ be a tree with root node $r$ and incidence matrix $\Omega$, if we denote the tail of $f$ by $F(t)_{x}\coloneqq\sum_{y}\mathbb{I}(y\in U(x))f(t)_{y}$, we can invert (4) via the tail distributions:
$\partial_{t}F(t)_{\lceil k\rceil}=v(t)_{k}g(t)_{k},$ (9)
@@
$W_{1}(f(0),f(1))=\inf_{f,g,v}\left\{\mathcal{I}_{2}(v,g):\partial_{t}=\Omega\cdot(vg))\right\}.$ (11)
@@
which transforms a sum over tails to a sum over edges. Squaring both sides and applying both Jensen’s and the Cauchy Schwarz inequalities bounds $W_{1}$ above by $V_{2}$.
@@
So, we have that $V_{2}\geq W_{1}$ too. However, Lemma 2, says that a constant speed solution achieves $W_{1}$, so choosing an arbitrary constant speed geodesic as in Definition 2 - see Appendix -B for an explicit example - we have that $V_{2}(f(0),f(1))=W_{1}(f(0),f(1))$ and we are done. ∎
In fact, the proofs for Lemma 2 and Theorem 1 work for $q\geq 1$ too, so we have a more general result that $V_{q}=W_{1}$, (see Appendix. -C) and we can choose $q$ arbitrarily.
@@
Consider Example 1 when p(t)=(1t)p+tqp(t)=(1-t)p+tq for $p>q>0$. We know that v(t)=n(qp)v(t)=n(q-p) and $g(t)=\text{Bin}(n-1,p(t))$ satisfy (4). We also have that $\mathcal{I}_{2}(v,g)=n(q-p)=W_{1}(f(0),f(1))$, so the original choice of $(f,v,g)$ is minimizing. This solution also contains a constant speed geodesic in the form of $f$.
@@ III-B The Reduced Formulation on a Graph
Given $f$, solving (5) for $(v,g)$ involves inverting $\Omega$. Fortunately there is a plethora of work on solving these systems, see [3, Section 2]. For an incidence matrix $\Omega$, we have $\text{Im}(\Omega)=\{x:\sum_{j}x_{j}=0\}$ and $\operatorname{rank}(\Omega)=|\mathcal{V}|-1$, see [3, Lemma 2.4]. We know $\partial_{t}f\in\text{Im}(\Omega)$ for all $t$, so we can always solve for the vector $vg$. Although, this system is not full rank. Removing an arbitrary row $r$ from $f$ and $\Omega$ gives
@@
and for shorthand we denote $f=\int\Omega\cdot(vg)$. Conversely, for any path $f(t)$ on the vertices, no matter which $(v,g)$ pair we choose as a solution to (4) we can reconstruct $f$ by using (12).
The lack of a unique inverse prevents us from obtaining $g$ in terms of $f$ and $v$ or $v$ in terms of $f$ and $g$. However we can use Remark 2, to limit our search for pairs $(v,g)$ living on the edges of $\mathcal{G}$.
@@
We have a new condition given by (13). So we should also investigate the structure of solutions, fortunately we can inherit most of the machinery from Remark 1, as we are still inverting an incidence matrix. We have that $(v,g)$ must satisfy
@@ III-C Benamou-Brenier on the Vertices of a graph
The multiple constant speed solutions explains why we cannot extend the argument of Section III-A to a graph. In fact not all constant speed solutions will minimize $V_{q}$, never mind integrate to $W_{1}$. However we can demonstrate that these solutions minimize $\mathcal{I}_{q}$ for a given $P$ and $\epsilon$.
@@
Given a graph $\mathcal{G}$ and $(P,\epsilon)$ a valid choice for inverting $\Omega$, define the Constant Speed Solution associated to $(P,\epsilon)$ to be
$v_{k}=\text{sign}(P\cdot\tilde{f}+\epsilon)_{k}|v|,\quad g_{k}=|P\cdot\tilde{f}+\epsilon|_{k}/|v|$ (15)
where $|v|=\sum_{k\in\mathcal{V}}|P\cdot\tilde{f}+\epsilon|_{k}$. Then the constant speed solution minimizes $\mathcal{I}_{q}(v,g)$ over the pairs $(v,g)$ associated to $(P,\epsilon)$, i.e pairs which satisfy (14).
@@
On a tree, we have one inverse (up to choice of root node) and Lemma 3 is equivalent to the constant speed solution minimising $\mathcal{I}_{q}(v,g)$, which is implied by Theorem 1.
@@
The following proposition demonstrates our motivation for considering time dependent pairs $(v,g)$, instead of fixed flows along edges, in that they induce a family of constant speed $W_{1}$, geodesics.
@@
Given a pair $(v,g)$ which achieves $W_{1}$, the induced path $f$ is a constant speed $W_{1}$ geodesic. Conversely, if $f$ is a constant speed geodesic, then there exists $(v,g)$ which achieves $W_{1}$ and induces $f$.
@@
Suppose that $(v,g)$ minimises $\mathcal{I}_{1}$. If it has constant speed, proceed, else substitute it for a constant speed solution as in Lemma 2. Then we define the path $f(t)=\int\Omega(vg)d\tau$ for $t\in[0,1]$. For an arbitrary subinterval $[s,t]$, the triple $(f,v,g)$ restricted to $[s,t]$ is still a solution to (5). What is unclear, is whether it is a minimising solution too. Suppose there exists $\tilde{v},\tilde{g}$ such that
$\int_{s}^{t}\sum_{k\in\mathcal{E}}\tilde{g}(\tau)_{k}|\tilde{v}(\tau)_{k}|d\tau<\int_{s}^{t}\sum_{k\in\mathcal{E}}g(\tau)_{k}|v(t)_{k}|dt.$
[…]
$(\hat{v},\hat{g})=\begin{cases}(\tilde{v},\tilde{g})&:\tau\in[s,t]\\ (v,g)&:\tau\not\in[s,t].\end{cases}$
But then by linearity of the integral we will have $\mathcal{I}_{1}(\tilde{v},\tilde{g})<\mathcal{I}_{1}(v,g)$ and hence $(v,g)$ is not a minimiser. So $(v,g)$ is a minimising solution between $[s,t]$ too and we have
$W_{1}(f(s),f(t))$ $=\int_{s}^{t}|v|d\tau$
$=(t-s)\int_{0}^{1}|v|d\tau=|t-s|W_{1}(f(0),f(1))$
and so $f$ is a constant speed geodesic.
Now suppose $f$ is a constant speed geodesic, we will use the result that the $W_{1}$ distance on a graph $\mathcal{G}$ is equal to $W_{1}$ across one of the spanning trees of $\mathcal{G}$, which we call $\mathcal{T}$. Then, we choose $(v,g)$ as the constant speed solution as in Lemma 2. This pair achieves $W_{1}$ on $\mathcal{T}$ and hence on $\mathcal{G}$. ∎
@@
IV Conclusion
We have extended the work of Hillion and Johnson (see [9, Lemma 3.6]) to demonstrate that the discrete Benamou-Brenier formulation of $W_{1}$ holds in generality on $\mathbb{Z}$, on trees, and eventually on graphs. The tree case lends itself particularly well, with an exact form for the minimizing triples in terms of tail distributions of a constant speed geodesics.
While the minimal Beckmann flow provides us with a convex interpolation (see Corollary 1), we see that on $\mathbb{Z}$ and even cycles we have minimizing triples $(f,v,g)$ with more exotic geodesics (see Examples 3,4 and Appendix -F). We formalise this characterisation of constant speed $W_{1}$ geodesics in Proposition 4.

### v2

@@ Abstract
We propose a discrete transport equation on graphs. We then derive a discrete analogue of the Benamou-Brenier formulation for Wasserstein-$1$ distance on a graph and as a result classify all constant speed $W_{1}$ geodesics on graphs.
@@ I Introduction
Let $\mu,\nu$ be a measures on $\mathcal{X}$ and $d$ a metric on $\mathcal{X}$, we define the set of couplings of $\mu$ and $\nu$ to be $\Pi(\mu,\nu):=\{\pi\in\mathcal{P}(\mathcal{X}^{2}):\pi(A\times\mathcal{X})=\mu(A),\pi(\mathcal{X}\times B)=\nu(B)\}$, and the Wasserstein $p$ distance by:
$W_{p}(\mu,\nu)^{p}=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathcal{X}^{2}}d(x,y)^{p}d\pi\right).$ (1)
It is known that $W_{p}$ is a metric on the space of (finite $p$’th moment) measures on $\mathcal{X}$, see [20, Definition 6.4]. The Wasserstein distance is popular in machine learning circles, namely as a loss function [2, 8, 17] and utilising its Riemannian structure for Wasserstein Gradient Flows [10, 15]. One notable development in optimal transport came via Benamou and Brenier [5], who re-parametrised the minimisation in terms of time-dependent distribution and velocity field pairs $(f_{t},v_{t})$.
@@
where $(f_{t},v_{t})$ are constrained by the transport equation:
$\partial_{t}f_{t}+\nabla\cdot(f_{t}v_{t})=0,$ (3)
@@
The infimum in (2) can be achieved by a constant speed $W_{p}$ geodesic $f_{t}$, and $v_{t}$ satisfying $v_{t}||=|\dot{f}|_{W_{p}}$, (see [5]).
On $\mathbb{R}^{d}$, the constant speed $W_{p}$ geodesics are the McCann displacement interpolations, see [14, Definition 1.1]. For discrete measures, say on $\{0,1...,n\}\subset\mathbb{R}$, the McCann interpolations will traverse $\mathbb{R}$, but not remain in the domain $\{0,1...,n\}$. This means to interpolate across discrete domains, considering them as embedded in a continuum will not suffice.
On a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, computing (1) is a $|\mathcal{V}|^{2}$-dimensional linear program. However, Beckmann [4] gave a time-independent velocity field formulation of $W_{1}$ which on a graph is $|\mathcal{E}|$-dimensional, (see [18, Section 6.3]). In fact the Benamou-Brenier formulation can be viewed as a generalisation of Beckmann’s work which introduces time parameters.
The main contribution to a discrete Benamou-Brenier analogue came from Erbar and Maas [13], who formulated an analogue of $W_{2}$ on discrete domains by considering a transport equation on graphs and a kinetic energy functional much like in Proposition 1. This metric is Riemannian and interplays well with discrete gradient flows, although importantly it is not the $W_{2}$ distance. This left the problem of a Benamou-Brenier interpretation of $W_{p}$ on graphs open.
Later, Hillion and Johnson [9] provided sufficient conditions for a true discrete Benamou-Brenier form of $W_{1}$ on $\mathbb{Z}$, which is the work we are building on. In this paper we offer our own discrete transport equation involving triples $(f,v,g)$, and derive a Benamou-Brenier formulation for $W_{1}$ on trees (Lemma 2, Theorem 1). We then generalise this result to graphs (Lemma 3, Theorem 2) and classify the constant speed $W_{1}$ geodesics via the solutions to our formulation (Proposition 4).
@@ II Notation and the Discrete Transport Equation
Consider the path of Binomial distributions $f(t)=\text{Bin}(n,p(t))$ on the set $\mathcal{V}=\{0,1,...,n\}$ where $p(t)$ varies with time. We show the case $n=5$ below:
@@
where $v(t)=np^{\prime}$ and $g(t)=\text{Bin}(n-1,p)$. Our key insight is that $\mathcal{V}$ can be considered as the vertices of a path graph $\mathcal{G}$, and that $g$ and $v$ exist on the edges of $\mathcal{G}$ as follows:
@@ II-A Matrices, Operators, Equations and Integrals
As shorthand for $\omega_{x,k}=1$ and $\omega_{k,x}=-1$ we will write $k\rightarrow x$ and $x\rightarrow k$ respectively.
@@
If we consider $\nabla_{1}$ as instead defined on edges (x,x1)(x,x-1) of $\mathbb{Z}$, it corresponds exactly to the $\nabla\cdot$ operator.
@@ II-B The Discrete Transport Equation
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with incidence matrix $\Omega$, we say the transport equation is satisfied if: for a path of distributions on vertices $f=f(t)_{x}$, a path of functions on edges $v=v(t)_{k}$ (called the velocity), and a path of distributions on edges $g=g(t)_{k}$, we have that for each $x\in\mathcal{V}$
$\partial_{t}f(t)_{x}+(\nabla\cdot v(t)g(t))_{x}\text{ for all }t\in[0,1].$ (4)
@@
and define $f$ to be $f(t)=(Z,Zs,Zs,Zs)$ for an arbitrary positive $s(t)$ and partition function $Z(t)=\frac{1}{1+3s}$. This is the stationary distribution of a Markov jump process $Q$ where $q_{x0}s(t)=q_{0x}$. Equation (4) states that
@@ III-A Benamou-Brenier on the Vertices of a Tree
It is worth noting that we have other special cases of $W_{1}$. For cyclic graphs, we have a another closed form expression, (see [19]), which utilises cuts much like Evans [7]. For a general graph, $W_{1}$ can be expressed as the minimum transport cost across possible spanning trees of $\mathcal{G}$, see [16].
@@
Let $\mathcal{G}$ be a tree with root node $r$ and incidence matrix $\Omega$, we can invert (4) via the tail distributions:
$\partial_{t}F(t)_{\lceil k\rceil}=v(t)_{k}g(t)_{k}\text{ for each }k\in\mathcal{E},$ (9)
@@
$W_{1}(f(0),f(1))=\inf_{f,g,v}\left\{\mathcal{I}_{2}(v,g):\partial_{t}f=\Omega\cdot(vg))\right\}.$ (11)
@@
which transforms a sum over tails to a sum over edges. Squaring both sides and applying both Jensen’s and the Cauchy–Schwarz inequalities bounds $W_{1}$ above by $V_{2}$.
@@
So, we have that $V_{2}\geq W_{1}$ too. However, Lemma 2 says that a constant speed solution achieves $W_{1}$, so choosing an arbitrary constant speed geodesic as in Definition 2 - see Appendix -B for an explicit example - we have that $V_{2}(f(0),f(1))=W_{1}(f(0),f(1))$ and we are done. ∎
In fact, the proofs for Lemma 2 and Theorem 1 work for $q\geq 1$ too, so we have a more general result that $V_{q}=W_{1}$, see Appendix -C , and we can choose $q$ arbitrarily.
@@
Consider the Binomial case (Example 1) when p(t)=(1t)p+tqp(t)=(1-t)p+tq for $p>q>0$. We know that v(t)=n(qp)v(t)=n(q-p) and $g(t)=\text{Bin}(n-1,p(t))$ satisfy (4). We also have that $\mathcal{I}_{2}(v,g)=n(q-p)=W_{1}(f(0),f(1))$, so the original choice of $(f,v,g)$ is minimizing. This solution also contains a constant speed geodesic in the form of $f$.
@@ III-B The Reduced Formulation on a Graph
Given $f$, solving (5) for $(v,g)$ involves inverting $\Omega$. Fortunately there is a plethora of work on solving these systems, see [3, Section 2]. For an incidence matrix $\Omega$, we have $\text{Im}(\Omega)=\{x:\sum_{j}x_{j}=0\}$ and $\operatorname{rank}(\Omega)=|\mathcal{V}|-1$, see [3, Lemma 2.4]. We know $\partial_{t}f\in\text{Im}(\Omega)$ for all $t$, so we can always solve for the vector $vg$. Although this system is not full rank, removing an arbitrary row $r$ from $f$ and $\Omega$ gives
@@
and for shorthand we denote $f=\int\Omega\cdot(vg)$. By construction, $(f,v,g)$ will satisfy (4). Conversely, for any path $f(t)$ on the vertices, no matter which $(v,g)$ pair we choose as a solution to (4) we can reconstruct $f$ by using (12).
The lack of a unique inverse of $\Omega$ prevents us from obtaining $g$ in terms of $f$ and $v$ or $v$ in terms of $f$ and $g$. However we can use Remark 2, to limit our search for pairs $(v,g)$ living on the edges of $\mathcal{G}$.
@@
We have a new condition given by (13), so we should also investigate the structure of solutions, fortunately we can inherit most of the machinery from Remark 1, as we are still inverting an incidence matrix. We have that $(v,g)$ must satisfy
@@ III-C Benamou-Brenier on the Vertices of a graph
The multiple possible inversions of $\Omega$ explain why we cannot extend the argument of Section III-A to a graph. In fact even constant speed solutions defined in Lemma 2 might not minimize $V_{q}$, never mind integrate to $W_{1}$. However we can demonstrate that constant speed solutions are still minimal.
@@
Given a graph $\mathcal{G}$, we say a pair $(v,g)$ which satisfies (13) is a Constant Speed Solution if $|v_{k}|=|v|$ for some function $|v|:[0,1]\rightarrow\mathbb{R}$. For any pair $(v,g)$ we have a constant speed solution $(v^{*},g^{*})$ such that $\mathcal{I}_{q}(v,g)\geq\mathcal{I}_{q}(v^{*},g^{*})$. Additionally, we can find another constant speed solution $(v^{**},g^{**})$ such that $|v^{**}_{k}|=|v^{**}|=c$ for $c\in\mathbb{R}$ and $\mathcal{I}_{q}(v^{*},g^{*})\geq\mathcal{I}_{q}(v^{**},g^{**})$.
@@
The key takeaway from Lemma 3 is that if we find any pair $(v,g)$ which achieves $W_{1}$, we can transform it via $(v^{*},g^{*})$ and $(v^{**},g^{**})$ to $W_{1}$ achieving pairs which have desirable properties. One of these properties will be key in the main result of Section IV. On a tree, we have one inverse (up to choice of root node) for Remark 3 and we know the constant speed solutions are explicitly defined via Lemma 2.
@@
IV Characterisation of Geodesics
On $\mathbb{R}^{d}$, we can achieve the minima of (2) by a pair $(f,v)$ including a constant speed geodesic $f$. We also know that constant speed $W_{p}$-geodesics between $f(0)$ and $f(1)$ are the McCann displacement interpolations – which are induced by an optimal couplings between $f(0)$ and $f(1)$.
In Examples 3, 4 and 5, we saw that the minimizing triple $(f,v,g)$ for (6) included a constant speed geodesic $f$ in much the same way. We also know in general that the path $f$ can be obtained from $(v,g)$ via (12) much like the McCann interpolations. We now establish our own characterisation of these geodesics.
@@ IV Characterisation of Geodesics
Given a pair $(v,g)$ which achieves $W_{1}$ in Proposition 3, and satisfies $|v_{k}|=|v|=c$, the induced path $f$ is a constant speed $W_{1}$ geodesic. Conversely, if $f$ is a constant speed geodesic, then there exists $(v,g)$ such that $|v_{k}|=|v|=c$ which achieves $W_{1}$ and induces $f$.
@@
Let $(v,g)$ be a constant speed solution that minimises $I_{1}$ and $|v_{k}|=|v|=c$, then induce the path $f(t)$. For an arbitrary subinterval $[s,t]$, the pair $(v,g)$ restricted to $[s,t]$ is still a solution to (3). It can also be shown by contradiction that $(v,g)$ restricted to $[s,t]$ also minimises $I_{1}$ across $[s,t]$, (see Appendix -F). Since $(v,g)$ is constant speed, we have
$W_{1}(f(s),f(t))=\int_{s}^{t}|v|d\tau$ $=(t-s)\int_{0}^{1}|v|d\tau$
$=(t-s)W_{1}(f(0),f(1))$
and so $f$ is a constant speed geodesic. Conversely, suppose $f$ is a constant speed geodesic, so $|\dot{f}|_{W_{1}}=W_{1}(f(0),f(1))$. We then use the remarkable result that (if we let $W^{\mathcal{H}}_{1}$ denote the Wasserstein distance across the graph $\mathcal{H}$) there exists some spanning tree $\mathcal{T}\subset\mathcal{G}$ such that $W^{\mathcal{T}}_{1}(f(0),f(1))=W^{\mathcal{G}}_{1}(f(0),f(1))$, (see [16]). We can choose $(v,g)$ as the constant speed solution on $\mathcal{E}_{\mathcal{T}}$ (as in Lemma 2), and $v=g=0$ for any $\mathcal{E}\backslash\mathcal{E}_{\mathcal{T}}$. This pair achieves $W_{1}^{\mathcal{T}}=W_{1}^{\mathcal{G}}$ and since $(f,v,g)$ satisfy the transport equation, $f$ is induced by $(v,g)$. ∎
@@
V Conclusion
We have extended the work of Hillion and Johnson (see [9, Lemma 3.6]) to demonstrate that the discrete Benamou-Brenier formulation of $W_{1}$ holds in generality on $\mathbb{Z}$, on trees, and eventually on graphs. The tree case lends itself particularly well, with an exact form for the minimizing triples in terms of tail distributions of constant speed , (see Lemma 1 and Theorem 1).
While the minimal Beckmann flow provides us with a convex interpolation (see Corollary 1), we see that on $\mathbb{Z}$ and even cycles we have minimizing triples $(f,v,g)$ with more exotic geodesics (see Examples 3, 4 and Example 6 in Appendix H). In fact the existence of Corollary 1 and Example 3 tell us that
$f_{1}(t)$ $=\text{Bin}(n,(1-t)p+tq)$
$f_{2}(t)$ $=(1-t)\text{Bin}(n,p)+t\text{Bin}(n,q)$
are two valid constant speed geodesics between $\text{Bin}(n,p)$ and $\text{Bin}(n,q)$. We formalise this characterisation of $W_{1}$ geodesics in Proposition 4.